OpenStudy (anonymous):

What is the product? 2/3*5/8*3/4 A.5/16 B.5/32 C.7/12 D.2/3

OpenStudy (mrnood):

\[\frac{ 2 }{ 3 }\times \frac{ 5 }{ 8 }\times \frac{ 3 }{ 4 }\] you simply multiply all the top numbers togehter then multiply all the bottom numbers together this gives you a single fraction - which you may need to reduce to its lowest form to get one of the nswers options

OpenStudy (mrnood):

OR look at the way my expression is laid out - you can simplify BEFORE multiplying: cancel the 3s and a cancel 2 against 4

OpenStudy (anonymous):

@Agl202

OpenStudy (mrnood):

@lamy why are you calling for help from someone else - ALL oyu nheed is above are you waiting fro someone to give you th eanswer?

OpenStudy (anonymous):

no i just dont know how to do it

OpenStudy (mrnood):

you know haw to cancel numbers from a fraction? there is a 3 on toip and bottom - so what does the expression look like when you cancel them...? also there is 2 on the top and 4 on the bottom - so what happens when you reduce that?

OpenStudy (mrnood):